what is the Domain of the function G(F(x))?
F(x) = x + 4, G(y) =
G(F(x)) = x+4

Answer by stanbon(75887)

what is the Domain of the function G(F(x))?
F(x) = x + 4, G(y) =
G(F(x)) = x+4
-------------------
G(F(x)) = G(x+4)
Comment: G is not defined so there is no answer.
Cheers,
Stan H.